# 61A Bot Report: AI Assistants in CS1 Save Students Homework Time and Reduce Demands on Staff. (Now What?)

### v2

@@ Abstract.
Chatbot interfaces for LLMs enable students to get immediate, interactive help on homework assignments, but even a thoughtfully-designed bot may not serve all pedagogical goals. In this paper, we report on the development and deployment of a GPT-4-based interactive homework assistant (“61A Bot”) for students in a large CS1 course; over 2000 students made over 100,000 requests of our bot across two semesters. Our assistant offers one-shot, contextual feedback, primarily through a low-friction “get feedback” prompt within the command-line “autograder” our students already run to test their code. Our Bot wraps student code in a custom prompt that supports our pedagogical goals and avoids providing solutions directly. We discuss our deployment and then analyze the impacts of our Bot on students, primarily through student-reported feedback and tracking of student homework progress. We find reductions in homework-related question rates in our course forum, as well as substantial reductions in homework completion time when our Bot is available. For students in the percentile, these reductions typically exceed 30 minutes per assignment, over 4 standard deviations faster than the mean in prior semesters. Finally, we conclude with a discussion of these observations, the potential impacts on student learning, as well as other potential costs and benefits of AI assistance in CS1.
@@ 1. Introduction
The recent wide availability of Large Language Models (LLMs) has given students in introductory Computer Science (CS) courses a tempting alternative to asking a human TA for help on programming assignments—and potentially waiting hours to receive it. However, while naively used LLMs do help students solve assigned problems, they typically do so by providing correct answers along with explanations, allowing students to avoid the process of developing solutions themselves and the learning associated with this process. A number of recent reports (Denny et al., 2023a, b; Finnie-Ansley et al., 2023; Hellas et al., 2023; Kazemitabaar et al., 2024; Liffiton et al., 2023; Wang et al., 2024) present more thoughtful approaches: new systems, also based on LLMs, geared towards offering guidance and assistance without providing direct solutions.
@@
Then, we examine our Bot’s impact on our CS1 course—a large (-student) Python-based course targeted at CS majors. Ultimately, we ask: What impact does this deployment have on our course, on students and on staff? First, we find that students assess a majority of Bot-provided hints as being helpful, in line with prior work. Second, through a retrospective observational study aimed at quantifying impacts, we find that requests for homework help to our course forum drop dramatically (75%) after our Bot is deployed, and that students spend substantially less time completing their homework—25-50% less time, often more than 30 minutes faster, a 4-8 standard deviation reduction from the mean completion time for the same assignments in prior semesters.
@@
Thus, our main contributions are:
A rich description of our low-friction CS1 assistant and its deployment in our course;
Findings of students’ subjective experiences based on student surveys over two semesters; and
A retrospective, observational analysis of student homework completion patterns, identifying quantitative impacts our bot had on staff demand and students’ homework experiences.
@@ 2. Background & Related Work
Generative models such as ChatGPT,111https://chat.openai.com/ OpenAI Codex (Chen et al., 2021), Amazon CodeWhisperer222https://aws.amazon.com/codewhisperer/, and GitHub Copilot333https://copilot.github.com/ offer promising opportunities for enriching the learning experience of students. These models have already been leveraged by educators in different areas of Computing education (Finnie-Ansley et al., 2022; Denny et al., 2023b; Hellas et al., 2023; Denny et al., 2023a), where they accelerate content generation and seem to be impacting the relevant skills students gain in introductory CS courses. Researchers have studied LLMs in areas such as generating code explanations (Leinonen et al., 2023a; Becker et al., 2023; MacNeil et al., 2023b; Liu et al., 2024a), providing personalized immediate feedback (Bassner et al., 2024), enhancing programming error messages (Leinonen et al., 2023b; Wang et al., 2024), generating discussion forum responses (Mitra et al., 2024; Liu et al., 2024b), and automatic creation of personalized programming exercises and tutorials (Sarsa et al., 2022; Yuan et al., 2023; Reeves et al., 2023) to enhance the comprehensiveness of course materials.
@@
Through deployments of LLMs as intelligent tutors, students can receive immediate, personalized support and guidance, which would ideally foster a deeper understanding of coding concepts and promote self-paced learning—just as with pre-LLM Intelligent Tutoring Systems (see (Crow et al., 2018) for a review, and (Suzuki et al., 2017) for an example). The ability of LLMs to generate tailored resources, such as new, personalized tutorials and newly-generated code examples, not only expands the available learning materials but also accommodates students’ varying learning preferences—though these generated materials are not always better (Pardos and Bhandari, 2023). Educators should integrate LLMs as complementary tools, striking a balance between automation and human interaction while emphasizing the development of critical problem-solving skills and responsible coding practices, ultimately serving students better in their CS education.
@@ 3. Design & Deployment
Our deployment focused on providing students help with homework problems in part to address frequent student feedback from prior terms about long wait times for TA support for homework problems. In particular, we chose to focus primarily on the kinds of debugging assistance our staff are often asked for.
Three concerns—hallucinations, students sharing personal information with a third party, and the harms from an unmonitored chat interaction—led us to deploy a one-shot “Get Help” interaction mode without an opportunity for “chat” follow-up. This meant that one valuable pedagogical tool—having students explain their understanding of the problem—would remain out of reach in this initial deployment.
@@
In addition to the steps above, the prompt also includes a per-problem instruction block, which we used for about 10% of problems, and more general instructions such as Do not give the student the answer or any code. and Limit your response to a sentence or two at most.
@@ 3.1. Course Details & Deployment
Our course covers most of the typical CS1 content, plus a few additional topics, and we report here on the CS1 portion of our course. These modules are taught using Python, to a student population made up predominantly of CS majors and intended majors ( majors/intended), with substantial prior preparation ( any prior CS course, including high school). These demographics are broadly consistent across the semesters we examine historically in this report, and variations do not correlate with the results described here. (Fall and Spring semester students do differ demographically due to how course requirements are structured at our institution, and so we treat those as separate populations for this work.) In our course, content is typically introduced in lecture, and then reinforced in three assignment types: first, “labs,” which include a mixture of notional machine reinforcement (e.g., tracing, “What would Python do?”) and traditional homework problems, completed in small groups without access to 61A Bot. These are followed by “homework,” which students complete individually with access to 61A Bot. Finally, students complete “projects” that bring together multiple concepts towards a single goal, over a few weeks, again without access to 61A Bot. In addition, students were prohibited by course policy from using ChatGPT or other similar AI-based systems for help across all assignment types—but we do not have the means to enforce this.
@@ 3.1.1. Using 61A Bot
Students primarily (93%) access our Bot through an autograder444 OK Client, https://github.com/okpy/ok-client in the command line which provides the result of running the student’s code on a set of test cases; if any fail, the student is then asked whether they would like to receive feedback on their code from 61A Bot (see Figure 1). If they do, the autograder collects the student’s code, any errors from executing test cases, and constructs a request from these.
Students using VS Code555Visual Studio Code, https://code.visualstudio.com (our course does not prescribe a specific editor) can install an extension to enable a “Get Help” button in the toolbar. When activated, the extension collects students’ code, makes a best guess of which homework problem the student is working on (as several problems often appear in a single file), and constructs a request.
In both cases, the collected code includes the student’s full assignment code file for that assignment, which includes code for multiple questions. Requests thus include: our prompt, along with any problem-specific notes; the text of the specific homework problem the student is requesting help for (typically one of 4-6 such problems within the assignment); up to 3 pairs of prior-code/Bot-response text (if the student has made prior requests); the student’s current code (for all problems within the assignment); and, finally, any error text from failed test cases. These requests are sent to a server run by our instructional staff, which passes it to GPT-4 and logs the request and GPT-4’s response for further analysis.
Students are informed when installing the software that, in using our assistant, all code they write will be sent to OpenAI via Microsoft Azure666https://ai.azure.com/, and that they should not include any content in their code files (e.g., comments) that they do not want to share.
@@ 3.1.2. Deployment Timeline
We piloted and continued development on 61A Bot throughout the academic year 2023-2024. In Fall 2023, we deployed an initial pilot of 61A Bot to a section of 400 students. This was followed by a full-scale deployment in week 10 (after the CS1 portion of our course) for the approximately 1400 students across both sections of the course. In conjunction with the wider deployment, we also enabled access through the autograder tool students could already run from the command line to validate their code against a set of test cases. In Spring 2024, all 900 enrolled students had access to both modalities of the bot from the start of the semester. Our qualitative analysis thus draws on both semesters, while our quantitative analysis is focused on the full deployment in Spring 2024 where we can more confidently compare like-for-like with a full-term deployment and historically comparable student populations.
@@ 4.2. Student Reception
In the Spring 2024 survey, we also asked two optional free-response questions: What did students like the most and least about learning with the bot? We include a representative response for each here:
Most liked: “What I loved about the bot is that it allowed me to get feedback when I didn’t have access to a tutor. Accordingly, instead of banging my head against the wall for hours, I was able to get feedback about what I was doing wrong and correct the mistakes. For me personally, I would have had a lot more success in this class if I would have had access to the bot for labs. (Labs on average took me about 2 and a half hours to complete, and sometimes longer if I didn’t have access to a tutor).” Least liked: “Sometimes the answers were slightly vague. Of course, the bot can’t simply spit out the answer, but sometimes it was frustrating how it would say ‘you’re on the right track, but there seems to be a conceptual misunderstanding with __’ — the explanations for the blank could be a bit jargon-filled and didn’t always directly help me resolve the misunderstanding due to imprecise language.”
@@ 4.3. Effects on Homework
To understand the effects of 61A Bot on students in our course, we compare student performance from our semester of full deployment, Spring 2024 (SP24), with performance from prior Spring semesters (see §3.1), going back through Spring 2021 (SP21)—comprising a total of 1,643,613 data points from 6,034 students. We chose this method of analysis in large part because our IRB does not allow differential access to tools in courses, and thus we could not run a randomized control trial.
Though 61A Bot is available to students for homework assignments, it is not available for lab assignments nor for projects (see §3.1). To the extent that performance differences on homework assignments between pre- and post-deployment semesters are inconsistent with performance differences on lab assignments and projects, some of this difference may be attributable to the use of the Bot. Though course content is substantially similar across the semesters we report on here, there is some variation in course staff, individual lectures, and specific problems within labs and homework assignments—where homework and lab problems have changed more than trivially, we have omitted them from our comparisons, examining only those that are identical in nature and sequencing within our course. (We confirm, too, that the specific assignments we report on here are representative of the full set of assignments, and we are aware of no other major changes in course content, delivery, or staffing occurred during this time.)
@@ 4.3.1. Data & Analytical Method
We calculate an approximate total completion time by summing the timestamp deltas between snapshots. To account for students completing the homework across multiple sessions, we ignore time deltas above a 60-minute threshold—a value chosen somewhat arbitrarily, but we confirm that the results we report here are robust to values in the range of 20 to 120 minutes.888We also conducted an independent analysis looking only at snapshot counts rather than summing time deltas. Those results were consistent with the findings we report here, as expected from (Piech et al., 2012); we report times here because we find the time metric to be more straightforward to reason about. For clarity, we aggregate individual repeated problems into “assignments,” e.g., “HW 1.” Occasionally, one assignment in a particular semester differs sufficiently from other semesters that we omit it entirely.999Like our choice of completion time vs. submission counts, we opted for this method of aggregation because we found it to be the clearest presentation of the effects we observed. We performed the analysis described in this section with individual problems on their own, as well as by comparing individual semesters with their prior (Spring) semesters and comparing across different subsets of problems and prior years; the results we report here are robust to all of the variations we tried, including many others that we do not report on here for space.
@@ 4.3.2. Results
The data indicate that the (Bot-available) homework completion times in Spring 2024 are between faster among students in the middle percentile of completion times. This represents a reduction in time of Standard Deviations (SD) compared with completion times from prior Spring semesters, a large effect.
In contrast, no-Bot lab completion times fall within 1 SD of the prior semesters’ mean for the first few labs. Most later labs lack enough consistency across semesters for a robust comparison, but lab 8 does offer a glimpse into one possible outcome: here, students in Spring 2024 overall took several standard deviations more time than the mean for students in prior semesters.
@@ 5. Discussion
Though we did not carefully examine (and don’t make claims about) student learning outcomes here, we have reason to believe that students are not performing dramatically worse after deployment. In particular, though student exam performance is sufficiently inconsistent (due to dramatic exam coverage differences from term to term) that we did not report on it here, it does not appear that Spring 2024 students performed substantially worse than in prior terms. Additionally, the differences in performance on assignments where the Bot was not available (labs and projects) were much smaller. We can’t say for certain, but we have not found much evidence for a major decrease in outcomes.
@@
On the costs side, with most homework help requests going through an LLM, our human TAs may struggle more to stay on top of common challenges among students, leading to a looser feedback loop. However, this challenge could be addressed with new tools that help TAs aggregate over automated help requests—but now with LLM-generated hints and student feedback on whether those hints are helpful. That is, of course, assuming CS1 courses do not elect to reduce staff in response.
Given the mode of access of 61A Bot, one likely change in student learning is a reduction in the ability to read and comprehend error traces. Now that our students have an easy way to get natural language feedback on failed test cases, they are much less incentivized to try to understand why a test case failed by reading the trace.
On the benefits side, instructors could use this extra student time to cover more material, such as effective debugging techniques or reading stack traces. Or, students could continue to simply spend less time on the course. Similarly, TAs could spend less time debugging straightforward homework errors and more time focused on other forms of support.

### v3

@@ Abstract.
LLM-based chatbots enable students to get immediate, interactive help on homework assignments, but even a thoughtfully-designed bot may not serve all pedagogical goals. We report here on the development and deployment of a GPT-4-based interactive homework assistant (“61A Bot”) for students in a large CS1 course; over 2000 students made over 100,000 requests of our Bot across two semesters. Our assistant offers one-shot, contextual feedback within the command-line “autograder” students use to test their code. Our Bot wraps student code in a custom prompt that supports our pedagogical goals and avoids providing solutions directly. Analyzing student feedback, questions, and autograder data, we find reductions in homework-related question rates in our course forum, as well as reductions in homework completion time when our Bot is available. For students in the percentile, reductions can exceed 30 minutes per assignment, up to 50% less time than students at the same percentile rank in prior semesters. Finally, we discuss these observations, potential impacts on student learning, and other potential costs and benefits of AI assistance in CS1.
@@ 1. Introduction
The recent wide availability of Large Language Models (LLMs) has given students in introductory Computer Science (CS) courses a tempting alternative to asking a human TA for help on programming assignments—and potentially waiting hours to receive it. For instructors of large courses, LLMs like ChatGPT appeal by offering 24/7 access to personalized assistance that will answer nearly any student question, though not always correctly, and with other risks.
But while naively used LLMs do help students solve assigned problems, they typically do so by providing correct answers along with explanations, allowing students to avoid the process of developing solutions themselves and the learning associated with this process. A number of recent reports (Denny et al., 2023a, b; Finnie-Ansley et al., 2023; Hellas et al., 2023; Kazemitabaar et al., 2024; Liffiton et al., 2023; Wang et al., 2024) present more thoughtful approaches: new systems, also based on LLMs, geared towards offering guidance and assistance without providing direct solutions.
@@
Then, we examine our Bot’s impact on our CS1 course, UC Berkeley’s CS 61A—a large (-student) Python-based course targeted at CS majors. Ultimately, we ask: What impact does this deployment have on our course, on students and on staff? First, we find that students assess a majority of Bot-provided hints as being helpful, in line with prior work. Second, through a retrospective observational study aimed at quantifying impacts, we find that requests for homework help to our course forum drop dramatically (75%) after our Bot is deployed, and that students spend substantially less time completing their homework—25-50% less time, often more than 30 minutes faster, for students at the same percentile rank in completion time across prior semesters.
@@
Our work makes three main contributions. First, it offers a rich description of our low-friction CS1 assistant and its deployment in our course. Second, we detail findings of students’ subjective experiences based on student surveys over two semesters. Third, we share results from a retrospective, observational analysis of student homework completion patterns, identifying quantitative impacts our Bot had on staff demand and students’ homework experiences.
@@ 2. Background & Related Work
Generative models such as ChatGPT,111https://chat.openai.com/ OpenAI Codex (Chen et al., 2021), Amazon CodeWhisperer,222https://aws.amazon.com/codewhisperer/ and GitHub Copilot333https://copilot.github.com/ offer promising opportunities for enriching the learning experience of students. These models have already been leveraged by educators in different areas of Computing education (Finnie-Ansley et al., 2022; Denny et al., 2023b; Hellas et al., 2023; Denny et al., 2023a), where they accelerate content generation and seem to be impacting the relevant skills students gain in introductory CS courses. Researchers have studied LLMs in areas such as generating code explanations (Leinonen et al., 2023a; Becker et al., 2023; MacNeil et al., 2023b; Liu et al., 2024a), providing personalized immediate feedback (Bassner et al., 2024), enhancing programming error messages (Leinonen et al., 2023b; Wang et al., 2024), generating discussion forum responses (Mitra et al., 2024; Liu et al., 2024b), and automatic creation of personalized programming exercises and tutorials (Sarsa et al., 2022; Yuan et al., 2023; Reeves et al., 2023) to enhance the comprehensiveness of course materials.
@@
Through deployments of LLMs as intelligent tutors, students can receive immediate, personalized support and guidance, which would ideally foster a deeper understanding of coding concepts and promote self-paced learning—just as with pre-LLM Intelligent Tutoring Systems (e.g., (Suzuki et al., 2017); see (Crow et al., 2018) for a review). The ability of LLMs to generate tailored resources, such as new, personalized tutorials and newly-generated code examples, not only expands the available learning materials but also accommodates students’ varying learning preferences—though these generated materials are not always better (Pardos and Bhandari, 2023). Educators should integrate LLMs as complementary tools, striking a balance between automation and human interaction while emphasizing the development of critical problem-solving skills and responsible coding practices, ultimately serving students better in their CS education.
@@
Our work here builds on these prior efforts in the design of 61A Bot, differing primarily in its integration mode (embedded within an autograder students must run anyway), and the lack of conversational interaction. Our evaluation validates prior findings on engagement and satisfaction, and additionally offers a unique perspective of the impact of our Bot on student assignment completion times compared with historical course baselines.
@@ 3. Design & Deployment
Our deployment focused on providing students help with homework problems in part to address frequent student feedback from prior terms about long wait times for TA support for homework problems. In particular, we chose to focus primarily on the kinds of debugging assistance our staff are often asked for. Three concerns—hallucinations, students sharing personal information with a third party, and the harms from an unmonitored chat interaction—led us to a one-shot “Get Help” design rather thanc“chat”. This meant eschewing the valuable pedagogical approach of having students explain their understanding of the problem.
@@
In addition to the steps above, the prompt also includes a per-problem instruction block, which we used for about 10% of problems, and more general instructions such as Do not give the student the answer or any code. and Limit your response to a sentence or two at most. Our full prompt, along with our server and VS Code extension, are open source and available online at https://github.com/zamfi/61a-bot.
@@ 3.1. Course Details & Deployment
Our course covers most of the typical CS1 content, plus a few additional topics, and we report here on the CS1 portion of our course. These modules are taught using Python, to a student population made up predominantly of CS majors and intended majors ( majors/intended), with substantial prior preparation ( any prior CS course, including high school). These demographics are broadly consistent across the semesters we examine historically in this report, and variations do not correlate with the results described here. (Typical Fall and typical Spring semester students differ demographically due to how course requirements are structured at our institution; we treat those as separate populations.) In our course, content is typically introduced in lecture, and then reinforced in three assignment types: first, “labs,” which include a mixture of notional machine reinforcement (e.g., tracing, “What would Python do?”) and traditional homework problems, completed in small groups without access to 61A Bot. These are followed by “homework,” which students complete individually with access to 61A Bot. Finally, students complete “projects” that bring together multiple concepts towards a single goal, over a few weeks, again without access to 61A Bot. In addition, students were prohibited by course policy from using ChatGPT or other similar AI-based systems for help across all assignment types—but we do not have the means to enforce this.
@@ 3.1.1. Using 61A Bot
Students primarily (93%) access our Bot through an autograder444 OK Client, https://github.com/okpy/ok-client in the command line which provides the result of running the student’s code on a set of test cases; if any fail, the student is then asked whether they would like to receive feedback on their code from 61A Bot (see Figure 1). If they do, the autograder collects the student’s code, any errors from executing test cases, and constructs a request from these. Alternatively, students choosing to use VS Code555Visual Studio Code, https://code.visualstudio.com can install an extension to enable a “Get Help” button in the toolbar that works similarly.
Requests thus include: our prompt, along with any problem-specific notes; the text of the specific homework problem the student is requesting help for (one of 4-6 problems within an assignment); up to 3 pairs of prior-code/Bot-response text (from any prior requests); the student’s current code; and, finally, any error text from failed test cases. These requests are sent to a server run by our instructional staff, which passes it to GPT-4 and logs the request and GPT-4’s response for further analysis.
Students are informed when installing the software that, in using our assistant, all code they write will be sent to OpenAI via Microsoft Azure,666https://ai.azure.com/ and that they should not include any content in their code files (e.g., comments) that they do not want to share.
@@ 3.1.2. Deployment Timeline
We piloted and continued development on 61A Bot throughout the academic year 2023-2024. In Fall 2023, we deployed an initial pilot of 61A Bot to a section of 400 students. This was followed by a full-scale deployment in week 10 (after the CS1 portion of our course) for the approximately 1400 students across both sections of the course. In conjunction with the wider deployment, we also enabled access through the autograder tool students could already run from the command line to validate their code against a set of test cases. In Spring 2024, all 900 enrolled students had access to both modalities of the Bot from the start of the semester. Our qualitative analysis thus draws on both semesters, while our quantitative analysis is focused on the full deployment in Spring 2024 where we can compare like-for-like with a full-term deployment and historically comparable student populations.
@@ 4.2. Student Reception
In the Spring 2024 survey, we also asked two optional free-response questions: What did students like the most and least about learning with the Bot? We include a representative response for each here: Most liked:
“What I loved about the bot is that it allowed me to get feedback when I didn’t have access to a tutor. Accordingly, instead of banging my head against the wall for hours, I was able to get feedback about what I was doing wrong and correct the mistakes. For me personally, I would have had a lot more success in this class if I would have had access to the bot for labs. (Labs on average took me about 2 and a half hours to complete, and sometimes longer if I didn’t have access to a tutor).”
Least liked:
“Sometimes the answers were slightly vague. Of course, the bot can’t simply spit out the answer, but sometimes it was frustrating how it would say ‘you’re on the right track, but there seems to be a conceptual misunderstanding with __’ — the explanations for the blank could be a bit jargon-filled and didn’t always directly help me resolve the misunderstanding due to imprecise language.”
@@ 4.3. Effects on Homework
To understand the effects of 61A Bot on students in our course, we compare student performance from our semester of full deployment, Spring 2024 (SP24), with performance from prior Spring semesters (see §3.1), going back through Spring 2021 (SP21)—comprising a total of 1,643,613 data points from 6,034 students. (Our IRB does not allow differential access to tools in courses, preventing a randomized control trial.)
Though 61A Bot is available to students for homework assignments, it is not available for lab assignments nor for projects (see §3.1). To the extent that performance differences on homework assignments between pre- and post-deployment semesters are inconsistent with performance differences on lab assignments and projects, some of this difference may be attributable to the use of the Bot. There is some variation in course staff, individual lectures, and specific problems within labs and homework assignments over time in our course; where problems have changed more than trivially in content or sequencing, we have omitted them from our comparisons. The specific assignments we report on here are otherwise representative of the full set of assignments.
∗ denotes differences significant at the level.
@@ 4.3.1. Data & Analytical Method
We calculate an approximate total completion time by summing the timestamp deltas between snapshots. To account for students completing the homework across multiple sessions, we ignore time deltas above a 60-minute threshold—a value chosen somewhat arbitrarily, but we confirm that the results we report here are robust to values in the range of 20 to 120 minutes.888An independent analysis of snapshot counts were consistent with the findings we report here, as expected from (Piech et al., 2012); we report times as we found them more straightforward to reason about. For clarity, we aggregate individual repeated problems into “assignments,” e.g., “HW 1.” Occasionally, one assignment in a particular semester differs sufficiently from other semesters that we omit it entirely.999We chose to aggregate in this way for clarity; our reported analyses are robust to aggregation method and problem selection, as well as many other factors elided here for space.
@@ 4.3.2. Results
The data indicate that the (Bot-available) homework completion times in Spring 2024 are between faster among students in the middle percentile of completion times, when compared to students at the same percentile rank (3(a)). There is some inherent variability in completion time from year to year for a given percentile rank (e.g., the percentile ranked student); the red bands in Figure 3 show this variability. Note that the Spring 2024 homework completion times within the percentile range were lower, at a given percentile rank, than the mean completion time at that rank across prior semesters—by a factor of the range of that prior variability across Spring semesters. In contrast, no-Bot lab completion times fall within the range of the prior semesters’ variability for the first few labs. Most later labs lack enough consistency across semesters for a robust comparison, but lab 8 does offer a glimpse into one possible outcome: here, students in Spring 2024 at a given rank took more time than students at the same rank in any prior semesters.
@@ 5. Discussion & Future Work
Though we did not carefully examine (and don’t make claims about) student learning outcomes, we have reason to believe that students are not performing much worse after deployment. In particular, student exam performance is sufficiently inconsistent (exam coverage differs from term to term) that we did not report on it here, but it does not appear that Spring 2024 students performed much worse than in prior terms. Additionally, the differences in performance on assignments where the Bot was not available (labs and projects) were much smaller. We can’t say for certain, but we have not found much evidence for a major decrease in outcomes.
@@
On the costs side, with most homework help requests going through an LLM, our human TAs may struggle more to stay on top of common challenges among students, leading to a looser feedback loop. However, this challenge could be addressed with new tools that help TAs aggregate over automated help requests—but now with LLM-generated hints and student feedback on whether those hints are helpful. That is, of course, assuming CS1 courses do not elect to reduce staff in response. Given the mode of access of 61A Bot, one likely change in student learning is a reduction in the ability to understand error traces. With an easy way to get natural language feedback on failed test cases, students are less incentivized to read traces, let alone understand them.
However, instructors could use this extra student time to cover more material, such as debugging techniques or reading traces. Or, students could continue to simply spend less time on the course. Similarly, TAs could spend less time debugging straightforward homework errors and more time focused on other forms of support.
@@
We ourselves plan to also explore new opportunities to extend 61A Bot in future work, in particular in tracking student progress over time, and in providing concept support though worked examples and links to our online textbook.